James Manufacturing has the following information available for July:

Actual Results Flexible Budget Variance Flexible Budget Sales Activity Variance Master Budget
Units 12,000 ? 3,000 U ?
Sales revenue ? $ 10,800 F ? ? ?
Less:
Variable manufacturing costs $ 90,500 $ 95,000 ? $ 127,000
Variable marketing and administrative ? $ 4,500 U ? $ 5,200 F $ 41,000
Contribution margin $ 60,000 ? ? $ 7,100 U ?

What was James’s actual sales revenue for July?

Multiple Choice

  • $190,800

  • $233,200

  • $180,000

  • $225,000
Answer #1
Variable marketing and administrative in flexible budget 35800 =41000-5200
Variable marketing and administrative in actual results 40300 =35800+4500
Contribution margin 60000
Add: Actual variable expenses 130800 =90500+40300
Actual sales revenue for July 190800
Option A $190,800 is correct